# Leveraging LLMs for User Stories in AI Systems: UStAI Dataset

### v1

@@ Abstract.
User stories, large language models, requirements quality, AI requirements, requirements elicitation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXbooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwayisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Requirements analysis
@@ 1. Introduction
  • Compile and curate a collection of user stories for AI systems dataset (UStAI) [^1] https://github.com/asmayamani/EthicsRequirementsData with $1260$ entries evaluated for quality attributes from the QUS framework and non-functional requirements, including implied ethical principles.
@@
Criteria Description Criteria type
Syntactic
Semantic
Pragmatic
Full-sentence A user story is a well-formed full sentence Individual
Unique Every user story is unique, duplicates are avoided Set
@@ 2.1. LLMs in requirements generation
UStAI (Ours) Synthetic using 3 LLMs: ChatGPT o1-mini, Gemini 1.5 flash, Llama 3.1 70b 1260 user stories about 48 ML systems Tabular data Yes Quality according to the QUS framework, High-level NFR, indicated ethics principles. 2024-2025 Quality evaluation, ethical issues identification, ambiguity detection, requirements categorization, conflict detection, and identification of equivalent requirements.
@@ 3. Study Design
User stories are a type of requirements artifact that describes software system features from the perspective of the users interacting with the system being developed (Inayat et al., 2015). Popularized by Mike Cohn (Cohn, 2004), user stories are typically written in the format “As a ⟨type of user⟩, I want ⟨goal⟩, so that ⟨some reason⟩” Known as the Connextra notation. User stories were chosen as the relation between the user, feature, reason, or end is essential to study the requirements for a human-centric approach.
Table 3. Sample of the generated user stories.
ID Title Machine Learning-Based Aggression Detection in Children with ADHD Using Sensor-Based Physical Activity Monitoring (Park et al., 2023)
A2US1O1 O1-mini As a parent, I want to receive real-time alerts when my child shows aggressive behavior, so I can intervene immediately.
A2US6O1 O1-mini As a child with ADHD, I want to wear a comfortable activity monitor, so my behavior is tracked without feeling uncomfortable.
A2US1Ge Gemini 1.5-flash As a parent, I want a device that can track my child’s physical activity and potentially identify aggressive episodes so that I can intervene and provide support.
A2US6Ge Gemini 1.5-flash As a child with a history of aggression, I want a wearable device that helps me and my parents monitor my activity levels and receive alerts when aggressive behavior might be escalating.
A2US1Ll Llama 3.1 70b As a researcher, I want to develop an objective method to track physical-aggressive incidents in children, so that I can identify early warning signs and provide timely interventions.
A2US6Ll Llama 3.1 70b As a developer, I want to design a wearable device (e.g., waist-worn activity monitor) that can track physical activity data in children and transmit it to a remote server for analysis, so that I can enable remote monitoring and detection of physical-aggressive incidents.
@@
Abstract selection $42$ abstracts were selected covering $26$ domains. $24$ of these domains were from the ontology developed by Fabris et al. (Fabris et al., 2022), which includes topics related to information and computer science, health, social science, economics and business, natural science, and other domains. In addition, we included abstracts related to autonomous vehicles and security, which are not a part of the Fabris et al. (Fabris et al., 2022) ontology, to ensure wider coverage. Twenty-six abstracts were from conference papers, while the remaining 14 were from Q1 or Q2 journals. The abstracts represent a system summary, as the construct of the abstract has the content needed in the user story in terms of background to identify the stakeholders ¡user¿, the methodology to identify the ¡goal¿, and the conclusion to identify the ¡reason¿. Abstracts also are available, accessible, and free. Moreover, the feasibility of many ML systems is explored through the development of ML components by ML researchers in academic institutes.
@@
We conducted a comprehensive evaluation by utilizing the Quality User Stories (QUS) framework proposed by Lucassen et al.(Lucassen et al., 2016). The QUS framework provides $13$ quality criteria for assessing the quality of user stories, grouped into three categories: syntactic, semantic, and pragmatic. The quality assessment taxonomy is shown in Table 1. Some of those criteria are related to the individual user stories, and some are related to the relation of the user story to other user stories in the same set. Lucassen et al. (Lucassen et al., 2016) also introduced the Automatic Quality User Story Artisan (AQUSA) tool to evaluate the quality of user stories, focusing on syntactic evaluation, uniqueness, and uniformity. We note here that we only evaluate semantic and pragmatic quality criteria if the user story is conceptually sound. This framework was also used partially to evaluate generated user stories in (Abed et al., 2024).
@@ 4. Results
A total of $1260$ user stories were generated. A sample of the generated user stories is in Table 3. The shortest user story generated was generated by Gemini-1.5 flash, while the longest was generated by Llama 3.1 70b. The average user story word count is  27 for Gemini and O1-mini and 31 for Llama generated user stories. The average user story word count for Llama generated user story is considered higher than user stories collected in (Dalpiaz, 2018) from different sources as the average lengths in  (Dalpiaz, 2018) ranged between 15-28 by source. More on the dataset statistics is in Table 4.
@@ 5.5. Dataset usage
  1. (1)
    Human-centric research related to Requirements in AI systems UStAI can be used to investigate ethical issues and concerns in various domains and as a seed dataset to elicit ethics-related requirements from the general population. Given that the dataset includes annotations for ethical principles for user stories that explicitly indicate ethical requirements, it can be used for AI-ethics-related research such as prioritization of ethics requirements, conflict resolution between ethical priorities for different stakeholders, and training models to detect ethical issues in software requirements documents. Moreover, as 55% of the user stories mention or imply at least one NFR, UStAI can be used to explore methods for automatically extracting and prioritizing NFRs in AI-based systems.
  2. (2)
    Benchmarking and Evaluation of LLMs for Requirements Engineering Researchers can use UStAI as a benchmark to compare LLMs in their ability to generate high-quality user stories based on the QUS framework with the quality of user stories generated by the three LLMs we investigated. Moreover, this dataset offers a basis for experimenting with various prompt engineering techniques for user story generation, leading to more effective ways of eliciting requirements from LLMs. In addition, UStAI can be used to investigate LLM’s ability to complete a set of requirements given a set of high-level requirements.
  3. (3)
    Training and Evaluation of NLP techniques: As UStAI is annotated with quality metrics, it can be used as training data for developing quality assessment tools using efficient and lightweight techniques. Also, it can be used as an evaluation dataset for existing methods and techniques for conflict detection, ambiguity detection, and other requirements smells. UStAI could be further annotated to be usable for other requirements engineering tasks.
@@ 5.6. Implications and future directions
However, current LLMs are not yet ready to fully support or replace human involvement in RE tasks. Several promising areas for future investigation remain. For example, researchers could explore the relevance of LLM-generated user stories in terms of scope, roles, and ethical considerations, as well as the alignment between ethics-related user stories and stakeholder values. Furthermore, agentic AI could be employed to enhance the completeness quality attribute or remove ambiguities especially for NFRs and ethics requirements. Another valuable direction is the empirical evaluation of LLMs to reduce the effort and cost associated with gathering requirements during the very early stages of RE. Also, investigating human-in-the-loop approaches to eliciting domain-specific requirements that require detailed expertise.
@@ 6. Conclusion
This work contributes to the growing body of knowledge on applied LLMs and the field of RE4AI. We proposed and evaluated generating user stories from abstract descriptions to build a synthetic AI user story dataset annotated with quality attributes, NFRs, and implied ethical principles. Our research demonstrated the potential of LLMs in producing high-quality user stories from abstract descriptions of AI components. It identified strengths and limitations in the syntactic, semantic, and pragmatic aspects and analyzed user stories for the presence of NFRs and ethical principles. The empirical evaluation concluded that Gemini 1.5 Flash outperformed other investigated LLMs by generating stories with fewer quality issues, greater emphasis on ethics, and a broader perspective, making it the preferred tool for early-stage requirements elicitation. We release the UStAI dataset of 1260 manually annotated user stories to the research community. Future work will explore automating quality evaluation and ethical alignment to advance tools designed to assist in eliciting ethical software requirements for AI systems.
@@ Acknowledgment
The authors would like to acknowledge the support received from Saudi Data and AI Authority (SDAIA) and King Fahd University of Petroleum and Minerals (KFUPM) under SDAIA-KFUPM Joint Research Center for Artificial Intelligence Grant no. JRC-AI-RG-12. The authors would also like to thank Eman Alofi, Nadeen Alamoudi, Mais Al-heraki, Linah Abouhajar for their contribution to the annotation and validation.

### v3

@@ Abstract.
User stories, large language models, quality requirements, requirements elicitation, requirements generation
ccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Requirements analysis
@@ 1. Introduction
  • Compile and curate a collection of user stories for AI systems dataset (UStAI) [^1] https://github.com/asmayamani/UStAI with $3000$ user stories including $1260$ user stories evaluated for quality attributes from the QUS framework and non-functional requirements, including implied ethical principles.
@@ 2.1. LLMs in requirements generation
UStAI (Ours) Synthetic using 3 LLMs: ChatGPT o1-mini, Gemini 1.5 flash, Llama 3.1 70b 3000 user stories from 100 ML systems, including 1260 evaluated for quality attributes and annotated with NFRs and ethics requirements Tabular data Yes Quality according to the QUS framework, High-level NFR, indicated ethics principles. 2024-2025 Quality evaluation, ethical issues identification, ambiguity detection, requirements categorization, conflict detection, and identification of equivalent requirements.
@@ 3. Study Design
Criteria Description Criteria type
Syntactic
Semantic
Pragmatic
Full-sentence A user story is a well-formed full sentence Individual
Unique Every user story is unique, duplicates are avoided Set
@@
User stories are a type of requirements artifact that describes software system features from the perspective of the users interacting with the system being developed (Inayat et al., 2015). Popularized by Mike Cohn (Cohn, 2004), user stories are typically written in the format “As a ⟨type of user⟩, I want ⟨goal⟩, so that ⟨some reason⟩,” known as the Connextra notation. User stories were chosen as the relation between the user, feature, reason, or end is essential to study the requirements for a human-centric approach.
Table 3. Annotated user stories for abstract with title Machine Learning-Based Aggression Detection in Children with ADHD Using Sensor-Based Physical Activity Monitoring (Park et al., 2023)(excerpt). Abbreviation key: Cjn = conjunctions, Min = minimal, CS = conceptually sound, PO = problem‑oriented, UA = unambiguous, CF = conflict‑free, Est = estimatable, Ind = independent, Uniq = unique, dup = isSemDuplicate, n.v.r = no_valid_role, SO = solution‑oriented, Gen = general.
ID US LLM Role User story NFR Ethics At. Mi. CS PO UA CF CFus Est. Indep. Uniq. Uniq.us M/E M/Eus
A2US1Ge US1 Gemini parent As a parent, I want a device that can track my child’s physical activity and potentially identify aggressive episodes so that I can intervene and provide support. N/A N/A Cjn Min CS PO UA CF Est Ind End 8
A2US2Ge US2 Gemini teacher As a teacher, I want a tool that can discreetly monitor student activity levels in the classroom and alert me to potential aggression so that I can de‑escalate situations before they escalate. N/A N/A Cjn Min CS PO UA Conf 9 Est No End 8
A2US9Ge US9 Gemini advocate As a data‐privacy advocate, I want assurances that any wearable device used for aggression detection collects data securely and protects children’s privacy. Privacy–Security Privacy / NM / Sec At Min CS PO Vag Src 2,3,4,5,8 Est Ind Uniq
A2US1Ll US1 Llama researcher As a researcher, I want to develop an objective method to track physical‑aggressive incidents in children, so that I can identify early warning signs and provide timely interventions. N/A Beneficence At Min n.v.r SO Gen CF No No Mean 10
A2US2Ll US2 Llama parent As a parent, I want to know if my child is experiencing aggressive episodes, so that I can provide support and guidance to help them manage their emotions and behaviors. N/A N/A At Min CS PO UA CF Est No dup 7 End
A2US7Ll US7 Llama parent As a parent, I want to receive alerts and notifications when my child experiences a physical‐aggressive incident, so that I can provide immediate support and guidance. Reliability N/A At Min CS PO UA CF Est No dup 2 End 2
A2US1o1 US1 o1‑mini parent As a parent, I want to receive real‑time alerts when my child shows aggressive behavior, so I can intervene immediately. Perf.–Reliab. N/A At Min CS PO UA CF Est No End
A2US2o1 US2 o1‑mini teacher As a teacher, I want daily reports on students’ aggression levels from wearable sensors, so I can implement effective classroom strategies. N/A N/A At Min CS PO UA Conf Est Ind diff End
A2US9o1 US9 o1‑mini officer As a privacy officer, I want all sensor data to be securely stored and access‑controlled, so children’s privacy is protected. Security–Privacy Privacy / Sec At Min CS PO UA Src Est Ind Uniq
@@
Abstract selection Initially, we generated a total of $3000$ user stories from $100$ scholarly abstracts covering diverse domains. However, to ensure a detailed, rigorous, and feasible evaluation, we selected randomly a representative subset consisting of $1260$ user stories from $42$ abstracts for in-depth analysis and annotation in this study covering $26$ domains. $24$ of these domains were from the ontology developed by Fabris et al. (Fabris et al., 2022), which includes topics related to information and computer science, health, social science, economics and business, natural science, and other domains. In addition, we included abstracts related to autonomous vehicles and security, which are not a part of the Fabris et al. (Fabris et al., 2022) ontology, to ensure wider coverage. Twenty-six abstracts were from conference papers, while the remaining 14 were from Q1 or Q2 journals. The abstracts represent a system summary, as the construct of the abstract has the content needed in the user story in terms of background to identify the stakeholders ¡user¿, the methodology to identify the ¡goal¿, and the conclusion to identify the ¡reason¿. Abstracts also are available, accessible, and free. Moreover, the feasibility of many ML systems is explored through the development of ML components by ML researchers in academic institutes.
@@
We conducted a comprehensive evaluation by utilizing the Quality User Stories (QUS) framework proposed by Lucassen et al.(Lucassen et al., 2016). The QUS framework provides $13$ quality criteria for assessing the quality of user stories, grouped into three categories: syntactic, semantic, and pragmatic. The quality assessment taxonomy is shown in Table 2. Some of those criteria are related to the individual user stories, and some are related to the relation of the user story to other user stories in the same set. We note here that we only evaluate semantic and pragmatic quality criteria if the user story is conceptually sound. Moreover, set-related quality criteria (e.g, conflict, uniqueness, .. etc) is evaluated with respect to the user stories from the same abstract and generated by the same LLM.The QUS framework is also applied in  (Abed et al., 2024) to evaluate user stories generated from stakeholder interviews. As no previous work was conducted on this type of work, we use a preliminary experiment conducted using ChatGPT-4 to generate user stories for ten scholarly abstracts as our baseline.
@@ 4. Results
A total of $3000$ user stories were generated, out of which $1260$ user stories from $42$ abstracts where selected randomly for analysis. The domain breakdown of the analyzed domains is in the appendix A. A sample of the generated user stories is in Table 3. The shortest user story generated was generated by Gemini-1.5 flash, while the longest was generated by Llama 3.1 70b. The average user story word count is  27 for Gemini and O1-mini and 31 for Llama generated user stories. The average user story word count for Llama generated user story is considered higher than user stories collected in (Dalpiaz, 2018) from different sources as the average lengths in  (Dalpiaz, 2018) ranged between 15-28 by source. More on the dataset statistics is in Table 4.
@@ 5. Discussion
Comparing to GPT‑4‑generated user stories, shown in Table 5, we can see the significant improvements when using the more advanced LLMs. The generated user stories from Gemini 1.5‑Flash, Llama 3.1 70 B, and o1‑mini are consistently more atomic, minimal, conceptually sound, and unambiguous than the earlier GPT‑4 baseline.However, when compared to a similar study in which practitioners crafted user stories based on interviews (Abed et al., 2024), it becomes evident that human-generated stories still set the benchmark for quality. These interview-derived stories outperformed all LLM-generated counterparts in atomicity, minimality, and conceptual soundness, with scores exceeding 88%. Nonetheless, estimatability and ambiguity remain persistent challenges for both LLM- and human-generated stories, as even the human-crafted set achieved only 77% in these aspects.
@@ 5.5. Dataset usage
1- Human-centric research related to Requirements in AI systems UStAI can be used to investigate ethical issues and concerns in various domains and as a seed dataset to elicit ethics-related requirements from the general population. Given that the dataset includes annotations for ethical principles for user stories that explicitly indicate ethical requirements, it can be used for AI-ethics-related research such as prioritization of ethics requirements, conflict resolution between ethical priorities for different stakeholders, and training models to detect ethical issues in software requirements documents. Moreover, as 55% of the user stories mention or imply at least one NFR, UStAI can be used to explore methods for automatically extracting and prioritizing NFRs in AI-based systems.
2- Benchmarking and Evaluation of LLMs for Requirements Engineering Researchers can use UStAI as a benchmark to compare LLMs in their ability to generate high-quality user stories based on the QUS framework with the quality of user stories generated by the three LLMs we investigated. Moreover, this dataset offers a basis for experimenting with various prompt engineering techniques for user story generation, leading to more effective ways of eliciting requirements from LLMs. In addition, UStAI can be used to investigate LLM’s ability to complete a set of requirements given a set of high-level requirements.
3- Training and Evaluation of NLP techniques: As UStAI is annotated with quality metrics, it can be used as training data for developing quality assessment tools using efficient and lightweight techniques. Also, it can be used as an evaluation dataset for existing methods and techniques for conflict detection, ambiguity detection, and other requirements smells. UStAI could be further annotated to be usable for other requirements engineering tasks.
@@ 5.6. Implications and future directions
However, current LLMs are not yet ready to fully support or replace human involvement in RE tasks, highlighting several promising areas for future research. For example, exploring alternative sources for generating user stories beyond scholarly abstracts, such as issue trackers or industrial reports. Additionally, further investigation is required to understand the performance variations among different LLMs in generating user stories, particularly regarding relevance in terms of scope, roles, and ethical considerations, as well as the alignment between ethics-related user stories and stakeholder values. Moreover, agentic AI could be employed to enhance completeness and reduce ambiguities, especially for non-functional and ethics-related requirements. Another valuable direction is the empirical evaluation of human-in-the-loop approaches for eliciting domain-specific requirements requiring detailed expertise. While this paper rigorously evaluates generated user stories using the QUS framework, future studies would benefit from validation involving real-world AI development teams or industry practitioners to assess practical applicability.
@@ 6. Conclusion
This work contributes to the growing body of knowledge on applied LLMs and the field of RE4AI. We proposed and evaluated generating user stories from abstract descriptions to build a synthetic AI user story dataset annotated with quality attributes, NFRs, and implied ethical principles. Our research demonstrated the potential of LLMs in producing high-quality user stories from abstract descriptions of AI components. It identified strengths and limitations in the syntactic, semantic, and pragmatic aspects and analyzed user stories for the presence of NFRs and ethical principles. The empirical evaluation concluded that Gemini 1.5 Flash outperformed other investigated LLMs by generating stories with fewer quality issues, greater emphasis on ethics, and a broader perspective, making it the preferred tool for early-stage requirements elicitation. We release the UStAI dataset of $1260$ manually annotated user stories and $1740$ generated non-annotated user stories to the research community. Future work will explore automating quality evaluation and ethical alignment to advance tools designed to assist in eliciting ethical software requirements for AI systems.
@@ Acknowledgment
The authors would like to acknowledge the support received from Saudi Data and AI Authority (SDAIA) and King Fahd University of Petroleum and Minerals (KFUPM) under SDAIA-KFUPM Joint Research Center for Artificial Intelligence Grant no. JRC-AI-RG-12. The authors would also like to thank Eman Alofi, Nadeen Alamoudi, Mais Al-heraki, and Linah Abouhajar for their contribution to the annotation and validation.
Appendix
Appendix A Domain frequency
The most frequently occurring domain in the UStAI is Health, appearing 7 times, followed closely by Information Systems with 6 occurrences and Autonomous Vehicles with 4 occurrences. Domains such as Law and Marketing appeared twice each. The remaining domains—including Social Media, Security, Biochemistry, Transportation, Economic, Pattern Recognition, Finance, Social Work, Library and Information Sciences, Movies, Sociology, Demography, Linguistics, Computer Networks, Signal Processing, Sports, Management Information Systems, Music, Social Network, Food, and Urban Studies—each occurred once.